A 750 g disk and a 760 g ring, both 15 cm in diameter, are rolling along a horizontal surface at 1.9 m/s when they encounter a 15 ∘ slope.

Part A) How far up the slope does the disk travel before rolling back down?

Part B) How far up the slope does the ring travel before rolling back down?
Answer #1

Tayo. Disk st Levi = 1.9 mts initial 15° energy conservation (PE)- + (KE). - (PE) + (KE)f ten (pei - o kes-o Kei = 1 mu? + 2due (kep = 3 mu? PE = mgh sin ig - h. xdisk The xsin iso (pfg - mg resin 150 put the all value in ean ④ ot 3 me = my resinisa 3 v ugsinis (Volig mis) (gog.sm/2 ddish 3 (1.9) 4X9.8X siniso Hisz 1.067 m AnsPaye __B) in Touls of the e serving 2/1.61 150 energy conservation (PE) + (kE) = (PE), + (KE) -- 0 (Pe= = KEP = 1 m ve ² + 1Page KEp = f mv 2 tl mv, KE— my;) PE = mgh Siniso = h h- & sinis ring ning PE = mgesin 15° ] put the all the value of in eaوهو حعا ولحنه - (9) gog sinis 23. الا)